OpenStudy (anonymous):

<1 and <2 are supplementary angles. m <1 is 4y + 7 and m <2 is 9y + 4. Find m<2. < = angle

OpenStudy (anonymous):

Love your profile pic

OpenStudy (anonymous):

Haha thanks :)

OpenStudy (anonymous):

Ok, supplementary angles always add up to 180 degress So m <1 + m <2 = 180 degrees 4y + 7 + 9y + 4 = 180 13y + 11 = 180 y=13 Plug y in for <2 m <2 = 9(13) + 4 m <2 = 121 God I wish I was doing what your doing in math right now.

OpenStudy (anonymous):

Great I did half of it right, thanks!

OpenStudy (anonymous):

Np, Ive encountered my fair share of unhelpful highschool teachers.

OpenStudy (anonymous):

Haha.